OpenStudy (anonymous):

how do you solve x-4 1 ---- = ------ x+6 7

OpenStudy (amistre64):

to start out, lets get everything out of fraction mode. multiply everything by 7, then multiply everything by (x+6)..what do we get for that?

OpenStudy (anonymous):

you mean taking the first part and multiplying it by 7 right then you would get 7x-4 and 7x+6 right?

OpenStudy (amistre64):

close, but we are not multiplying by 7/7...that would not help... 7(x-4) 1 (7) ----- = ------ what does this equal? (x+6) 7

OpenStudy (anonymous):

sorry it is 7x-28 and 7 which makes the second half a whole number of 1 right

OpenStudy (amistre64):

if I understand you correctly, then yes; we get: 7x-28 ----- = 1 (x+6) now we multiply both sides by (x+6): (x+6) (7x-28) ------------ = 1(x+6) what does this do for us? (x+6)

OpenStudy (anonymous):

7x^2-168 = x+6

OpenStudy (amistre64):

close.... but lets see if we can see what happens here: whenever a number is divided by itself it equals 1: 4/4 = 1 7/7 = 1 (x+6)/(x+6) = 1 (x+6) (7x-28) ----- ------- = 1(x+6) means: (x+6) 1(7x-28) = 1(x+6) right?

OpenStudy (anonymous):

7x-28= x+6

OpenStudy (amistre64):

very good.. now do you know how to solve it from here?

OpenStudy (anonymous):

then you subtract the x from x+6

OpenStudy (amistre64):

yes... and whatever you do to one side, you have to do the same thing to the other... so lets subtract x from both sides: 7x-x -28 = x-x +6 6x - 28 = 0+6 6x -28 = 6 right? now what?

OpenStudy (anonymous):

add 28 to the 6 then you get 6x=34 then divide both sides by 6 and you get 5.6

OpenStudy (amistre64):

very good :) but I dont think.... the exact amount would be 34/6 which can be reduced to 17/3. 17/3 = 5.666666666666666........ the (6)s never end.

OpenStudy (amistre64):

we can keep it as 17/3, or estimate its value to 5.6; or maybe 5.7

OpenStudy (anonymous):

awesome thanks for the help it was nice to have some one to work the problem with me

OpenStudy (amistre64):

youre welcome :)